Use Dijkstra’s algorithm to find the shortest path spanning tree for the following weighted 

directed graph with vertices A, B, C, D, and E given. Consider the starting vertex as E.




1
Expert's answer
2021-06-03T08:23:50-0400

For example:



shortest path to D: E - D, cost=2

shortest path to C: E - C, cost=3

shortest path to B: E - C - B, cost=4

shortest path to A: E - C - A, cost=6


The minimum spanning tree: